OTOnits 10, 10 years ago, the City ofMelrose issued $4,000,000 of9% coupon, 15-year, annual payment, tax-exempt muni bonds. The bonds had 10 years of call protection, but now the bonds can be called if the city chooses to do so. The call premium would be 5% of the face amount. New 5-year, 7%, annual payment bonds can be sold at par, but flotation costs on this issue would be 3% of the amount of bonds sold. What is the net present value of the refunding? Note that cities pay no income taxes, hence taxes are not relevant.
Answer #1

Current Annual coupons = 4,000,000 x 9% = 360,000

New annual coupon if bonds are refunded = 4,000,000 x 7% = 280,000

Annual saving = 80,000

Now in order to saving that, upfront cost include

Call premium = 5% x 4,000,000 = 200,000

Floating cost = 4,000,000 x 3% = 120,000

Let's calculate present value of annual saving using PV function

N = 5, PMT = 80,000, FV = 0, I/Y = 7%

=> Compute PV = $328,015.79

NPV = 328,015.79 - 120,000 - 200,000 = $8,015.79